The length of a rectangle is 7 centimeters less than its width. What are the
dimensions of the rectangle if its area is 228 square centimeters? The length of the rectangle is _____cm, The width of the rectangles is __cm.

Answer by jorel1380(3719)

Let l be the length of the rectangle. Then its width is l+7. So:
l(l+7)=228
lē+7l-228=0
(l+ 19)(l-12)=0
l=12 or -19
Using the positive value, we get 12cm as the length, and 19cm as the width. ☺☺☺☺